OpenStudy (anonymous):

Evaluate a-[b-(a-b)] if a=5 and b=3

OpenStudy (anonymous):

4

OpenStudy (anonymous):

u have to work out the inner parenthesses first inside the brackets 5-3=2 then 3-2=1 then5-1=4

OpenStudy (anonymous):

thx